# Are Self-Attentions Effective for Time Series Forecasting?

### v1

@@ Abstract
Time series forecasting is crucial for applications across multiple domains and various scenarios. Although Transformer models have dramatically shifted the landscape of forecasting, their effectiveness remains debated. Recent findings have indicated that simpler linear models might outperform complex Transformer-based approaches, highlighting the potential for more streamlined architectures. In this paper, we shift focus from the overall architecture of the Transformer to the effectiveness of self-attentions for time series forecasting. To this end, we introduce a new architecture, Cross-Attention-only Time Series transformer (CATS), that rethinks the traditional Transformer framework by eliminating self-attention and leveraging cross-attention mechanisms instead. By establishing future horizon-dependent parameters as queries and enhanced parameter sharing, our model not only improves long-term forecasting accuracy but also reduces the number of parameters and memory usage. Extensive experiment across various datasets demonstrates that our model achieves superior performance with the lowest mean squared error and uses fewer parameters compared to existing models.
@@ 1 Introduction
Time series forecasting plays a critical role within the machine learning society, given its applications ranging from financial forecasting to medical diagnostics. To improve the accuracy of predictions, researchers have extensively explored and developed various models. These range from traditional statistical methods to modern deep learning techniques. Most notably, Transformer Vaswani et al. (2017) has brought about a paradigm shift in time series forecasting, resulting in numerous high-performance models, such as Informer Zhou et al. (2021), Autoformer Wu et al. (2021), Pyraformer Liu et al. (2021), FEDformer Zhou et al. (2022b), and Crossformer Zhang and Yan (2022). This line of work establishes new benchmarks for high performance in time series forecasting.
However, recently, Zeng et al. (2023) have raised questions about the effectiveness of Transformer-based time series forecasting models especially for long term time series forecasting. Specifically, their experiments demonstrated that simple linear models could outperform these Transformer-based approaches, thereby opening new avenues for research into simpler architectural frameworks. Indeed, the following studies Liu et al. (2022); Ekambaram et al. (2023) have further validated that these linear models can be enhanced by incorporating additional features.
Despite these developments, the effectiveness of Transformer architecture in time series forecasting remains a subject of debate. Nie et al. (2023) introduced an encoder-only Transformer that utilizes patching rather than point-wise input tokens, which exhibited improved performance compared to linear models. Zeng et al. (2023) also highlighted potential shortcomings in simpler linear networks, such as their inability to capture temporal dynamics at change points Van den Burg and Williams (2020) compared to Transformers. Consequently, while a streamlined architecture may be beneficial, it is imperative to critically evaluate which elements of the Transformer are necessary and which are not for time series modeling.
In light of these considerations, our study shifts focus from the overall architecture of the Transformer to a more specific question: Are self-attentions effective for time series forecasting? While this question is also noted in Zeng et al. (2023), their analysis was limited to substituting attention layers with linear layers, leaving substantial room for potential model design when focusing on Transformers. Furthermore, the issue of temporal information loss (i.e., permutation-invariant and anti-order characteristics of self-attention) is predominantly caused by the use of self-attention rather than the Transformer architecture itself. Therefore, we aim to resolve the issues of self-attention and therefore propose a new forecasting architecture that achieves higher performance with a more efficient structure.
@@ 2 Related Work
Transformer models Vaswani et al. (2017) have shown effective in various domains Dosovitskiy et al. (2021); Devlin et al. (2019); Radford et al. (2018), with a novel encoder-decoder structure with self-attention, masked self-attention, and cross-attention. The self-attention mechanism is a key component for extracting semantic correlations between paired elements, even with identical input elements; however, autoregressive inference with self-attention requires quadratic time and memory complexity. Therefore, Informer Zhou et al. (2021) proposed directly predicting multi-steps, and a line of work, such as Autoformer Wu et al. (2021), FEDformer Zhou et al. (2022b), and Pyraformer Liu et al. (2021), investigated the complexity issue in time series transformers. Simultaneously, unique properties of time series, such as stationarity Liu et al. (2022), decomposition Wu et al. (2021), frequency features Zhou et al. (2022b), or cross-dimensional properties Zhang and Yan (2022) were employed to modify the attention layer for forecasting tasks. Recently, researchers have investigated the essential architecture in transformers to capture long-term dependencies. PatchTST Nie et al. (2023) became a de-facto standard Transformer model by patching the time series input in a channel-independence manner, which is widely used in following transformer forecasting models Liu et al. (2024); Goswami et al. (2024). On the other hand, Das et al. (2024) emphasized the importance of decoder-only forecasting models, while they focused on zero-shot using pre-trained language models. However, none of them have investigated the importance of cross-attention for time series forecasting.
@@
Fixed temporal order in time series is the distinct property of time series, in contrast to the language domain where semantic information does not heavily depend on the word ordering Devlin et al. (2019). Thus, some researchers have used learnable positional encoding in transformer models to embed time-dependent properties Li et al. (2019); Wu et al. (2021). However, Zeng et al. (2023) first argued that self-attention is not suitable for time series due to its permutation invariant and anti-order properties. While they focus on building complex representations, they are inefficient in maintaining the original context of historical and future values. They rather proposed linear models without any embedding layer and demonstrated that it can achieve better performance than transformer models, particularly showing robust performance to long input sequences. Recent linear time series models outperformed previous transformer models with simple architectures by focusing on pre-processing and frequency-based properties Li et al. (2023); Chen et al. (2023); Wang et al. (2024). On the other hand, Woo et al. (2023) investigated the new line of works of time-index models, which try to model the underlying dynamics with given time stamps. These related works imply that preserving the order of time series sequences plays a crucial role in time series forecasting.
3 Proposed Methodology
@@ 3.1 Problem Definition and Notations
In traditional time series transformers, we feed the historical multivariate time series $\bm{X}$ to embedding layers, resulting in the historical embedding $\bm{H}_{L}\in\mathbb{R}^{D\times L}$. Here, $D$ is the embedding size. Note that, in channel-independence manners, the multivariate input is considered to separate univariate time series ${\mathbf{x}}\in\mathbb{R}^{1\times L}$. With patching Nie et al. (2023), univariate time series ${\mathbf{x}}$ transforms into patches ${\mathbf{p}}=\texttt{Patch}({\mathbf{x}})\in\mathbb{R}^{P\times N_{L}}$ where $P$ is the size of each patch and $N_{L}$ is the number of input patches. Similar to non-patching situations, patches are fed to embedding layers $\bm{H}_{L}=\texttt{Embedding}({\mathbf{p}})\in\mathbb{R}^{D\times N_{L}}$.
Algorithm 1 Transformer Algorithm Vaswani et al. (2017)
[…]
In Fig. 2, we illustrate the existing time series forecasting architectures and their key components. Transformer-based models, as shown in Fig. 2a, exhibit the most complex architecture. These models leverage Self-Attention (SA), Masked Self-Attention (MSA), and Cross-Attention (CA) to generate predictions. The detailed algorithm is formalized in Algorithm 1, where MLP represents the fully connected layer and LN represents LayerNorm. Note that, as input tokens $\bm{H}_{T}$ for cross-attention, positional embedding is often used Zhou et al. (2021); Wu et al. (2021). The output from the cross-attention, $\bm{Z}^{(\texttt{Dec})}_{T}\in\mathbb{R}^{D\times N_{T}}$, is subsequently used to produce the final prediction $\bm{\hat{X}}$ through additional layers. Particularly, if we remove the decoder structure from this transformer architecture as shown in Figure 2b, we call them encoder-only models Liu et al. (2024); Nie et al. (2023).
In contrast, as shown in Fig. 2c, simplified linear models, such as DLinear Zeng et al. (2023), remove all transformer-based components and rebuild direct linear connections between input and output data. By removing these transformer-based components, these models avoid temporal information loss problems Zeng et al. (2023) and the heavy computational load that increases quadratically with the input size $L$, i.e., $\mathcal{O}(L^{2})$. However, we emphasize that this temporal information loss and computational overhead, coupled with self-attention rather than the transformer architecture itself.
@@ 3.2 Model Structure
We propose a model that not only preserves the temporal information of time series data similar to linear models but also utilizes the structural advantages of the transformer architecture. As illustrated in Figure 2d, cross-attention Transformer networks can maintain the periodic properties of time series, except for self-attention, which has permutation-invariant and anti-order characteristics. While Zeng et al. (2023) replaced all attention layers with linear layers, we argue that this approach does not fully address the underlying issue; namely, the potential of the transformer architecture itself, excluding self-attention, has been overlooked. Therefore, we introduce a novel approach that prioritizes cross-attention without self-attention, leveraging advanced architectural designs of time series transformers, which linear models cannot utilize, such as patching Nie et al. (2023).
Our proposed architecture consists of three key components: (A) Cross-Attention with Future as Query, (B) Parameter Sharing across Horizons, and (C) Query-Adaptive Masking. Fig. 3 illustrates how our model modifies the traditional transformer structure by removing self-attention and instead incorporating cross-attention, utilizing future data as the query. Additionally, we simplify the architecture by parameter sharing across forecasting horizons. We further enhance the performance through query-adaptive masking. The following paragraphs provide detailed descriptions of each component.
@@
Similar to self-attention, the cross-attention mechanism employs three elements: key, query, and value. The distinctive feature of cross-attention is that the query originates from a different source than the key or value. Generally, the query component aims to identify the most relevant information among the keys and uses it to extract crucial data from the values Chen et al. (2021); Zhang et al. (2023). In the realm of time series forecasting, where predictions are often made for a specific target horizon—such as forecasting 10 steps ahead. Therefore, within this concept of forecasting, we argue that each future horizon should be regarded as a question, i.e., an independent query.
To implement this, we establish horizon-dependent parameters as learnable queries. As shown in Fig. 3, we begin by creating parameters for the specified forecasting horizon. For each of these virtualized parameters, we assign a fixed number of parameters to represent the corresponding horizon as learnable queries ${\mathbf{q}}$. For example, ${\mathbf{q}}_{i}$ is a horizon-dependent query at $L+i$. When patching is applied, these queries are then processed independently: each learnable query ${\mathbf{q}}\in\mathbb{R}^{P}$ is first fed into the embedding layer, and then fed into the multi-head attention with the embedded input time series patches as the key and value.
@@
Transformer Liu et al. (2024) $\mathcal{O}(L^{2})$ $\mathcal{O}(T(T+L))$ 10.4ms Informer Zhou et al. (2021) $\mathcal{O}(L\log L)$ $\mathcal{O}(T(T+\log L))$ 13.5ms
Autoformer Wu et al. (2021) $\mathcal{O}(L\log L)$ $\mathcal{O}\left(\left(L/2+H\right)\log\left(L/2+T\right)\right)$ 24.1ms Pyraformer Liu et al. (2021) $\mathcal{O}(L)$ $\mathcal{O}(T(T+L))$ 11.2ms
FEDformer Zhou et al. (2022b) $\mathcal{O}(L)$ $\mathcal{O}\left(L/2+H\right)$ 69.3ms Crossformer Zhang and Yan (2022) $\mathcal{O}\left({M}L^{2}/{P^{2}}\right)$ $\mathcal{O}\left({M}T(T+L)/{P^{2}}\right)$ 30.6ms
PatchTST Nie et al. (2023) $\mathcal{O}\left({L^{2}}/{P^{2}}\right)$ 7.6ms CATS (Ours) - $\mathcal{O}\left({LT}/{P^{2}}\right)$ 7.0ms
@@
This independent forecasting mechanism offers a notable advantage: a higher level of parameter sharing. As demonstrated in Nie et al. (2023), significant reductions in the required number of parameters can be achieved in time series forecasting through parameter sharing between inputs or patches, enhancing computational efficiency. Regarding this, we propose parameter sharing across all possible layers — the embedding layer, multi-head attention, and projection layer — for every horizon-dependent query ${\mathbf{q}}$. In other words, all horizon queries ${\mathbf{q}}_{1},\dots,{\mathbf{q}}_{T}$ or ${\mathbf{q}}_{1},\dots,{\mathbf{q}}_{{N_{T}}}$ share the same embedding layer used for the input time series ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{L}$ or patches ${\mathbf{p}}_{1},\dots,{\mathbf{p}}_{N_{L}}$ before proceeding to the cross-attention layer, respectively. Furthermore, to maximize the parameter sharing, we also propose cross-dimension sharing that uses the same query parameters for all dimensions.
For the multi-head attention and projection layers, we apply the same algorithm across horizons. Notably, unlike the approach in PatchTST Nie et al. (2023), we also share the projection layer for each prediction. Specifically, PatchTST employs a fully connected layer as the projection layer for the concatenated outputs $\bm{Z}^{(\texttt{Dec})}_{T}$. Thus, its number of parameters becomes $(D\times N_{L})\times T$. However, our model shares the same projection layer for each prediction. The number of parameters becomes $D\times P$, which is not proportionally increasing to $T$. This approach significantly reduces time complexity during both the training and inference phases. The details are presented in Appendix.
@@
Therefore, to ensure the model focuses on each horizon-dependent query ${\mathbf{q}}$, we introduce a new technique that masks the attention outputs. As illustrated in the right-bottom figure of Fig. 3, for each horizon, we apply a mask to the direct connection from Multi-Head Attention to LayerNorm with a probability $p$. This mask prevents access to the input time series information, resulting in only the query to influence future value predictions. This selective disconnection, rather than the application of dropout in the residual connections, helps the layers to concentrate more effectively on the forecasting queries. We note that this approach can be related to stochastic depth in residual networks Huang et al. (2016). The stochastic depth technique has proven effective across various tasks, such as vision tasks Strudel et al. (2021); Yang et al. (2022). To the best of our knowledge, this is the first application of stochastic depth in Transformers for time series forecasting. A detailed analysis of query-adaptive masking can be found in Appendix.

### v3

@@ Abstract
Time series forecasting is crucial for applications across multiple domains and various scenarios. Although Transformers have dramatically advanced the landscape of forecasting, their effectiveness remains debated. Recent findings have indicated that simpler linear models might outperform complex Transformer-based approaches, highlighting the potential for more streamlined architectures. In this paper, we shift the focus from evaluating the overall Transformer architecture to specifically examining the effectiveness of self-attention for time series forecasting. To this end, we introduce a new architecture, Cross-Attention-only Time Series transformer (CATS), that rethinks the traditional transformer framework by eliminating self-attention and leveraging cross-attention mechanisms instead. By establishing future horizon-dependent parameters as queries and enhanced parameter sharing, our model not only improves long-term forecasting accuracy but also reduces the number of parameters and memory usage. Extensive experiment across various datasets demonstrates that our model achieves superior performance with the lowest mean squared error and uses fewer parameters compared to existing models. The implementation of our model is available at: https://github.com/dongbeank/CATS.
11footnotetext: Corresponding authors
@@ 1 Introduction
Time series forecasting plays a critical role within the machine learning society, given its applications ranging from financial forecasting to medical diagnostics. To improve the accuracy of predictions, researchers have extensively explored and developed various models. These range from traditional statistical methods to modern deep learning techniques. Most notably, Transformer [19] has brought about a paradigm shift in time series forecasting, resulting in numerous high-performance models, such as Informer [29], Autoformer [23], Pyraformer [11], FEDformer [31], and Crossformer [28]. This line of work establishes new benchmarks for high performance in time series forecasting.
However, Zeng et al. [26] have raised questions about the effectiveness of Transformer-based time series forecasting models especially for long term time series forecasting. Specifically, their experiments demonstrated that simple linear models could outperform these Transformer-based approaches, thereby opening new avenues for research into simpler architectural frameworks. Indeed, the following studies [12, 6] have further validated that these linear models can be enhanced by incorporating additional features.
Despite these developments, the effectiveness of each components in Transformer architecture in time series forecasting remains a subject of debate. Nie et al. [14] introduced an encoder-only Transformer that utilizes patching rather than point-wise input tokens, which exhibited improved performance compared to linear models. Zeng et al. [26] also highlighted potential shortcomings in simpler linear networks, such as their inability to capture temporal dynamics at change points [18] compared to Transformers. Consequently, while a streamlined architecture may be beneficial, it is imperative to critically evaluate which elements of the Transformer are necessary and which are not for time series modeling.
In light of these considerations, our study shifts focus from the overall architecture of the Transformer to a more specific question: Are self-attentions effective for time series forecasting? While this question is also noted in [26], their analysis was limited to substituting attention layers with linear layers, leaving substantial room for potential model design when focusing on Transformers. Furthermore, the issue of temporal information loss (i.e., permutation-invariant and anti-order characteristics of self-attention) is predominantly caused by the use of self-attention rather than the Transformer architecture itself. Therefore, we aim to resolve the issues of self-attention and therefore propose a new forecasting architecture that achieves higher performance with a more efficient structure.
@@ 2 Related Work
Transformer models [19] have shown effective in various domains [5, 4, 15], with a novel encoder-decoder structure with self-attention, masked self-attention, and cross-attention. The self-attention mechanism is a key component for extracting semantic correlations between paired elements, even with identical input elements; however, autoregressive inference with self-attention requires quadratic time and memory complexity. Therefore, Informer [29] proposed directly predicting multi-steps, and a line of work, such as Autoformer [23], FEDformer [31], and Pyraformer [11], investigated the complexity issue in time series transformers. Simultaneously, unique properties of time series, such as stationarity [12], decomposition [23], frequency features [31], or cross-dimensional properties [28] were employed to modify the attention layer for forecasting tasks. Recently, researchers have investigated the essential architecture in Transformers to capture long-term dependencies. PatchTST [14] became a de-facto standard Transformer model by patching the time series input in a channel-independence manner, which is widely used in following Transformer-based forecasting models [13, 7]. On the other hand, Das et al. [3] emphasized the importance of decoder-only forecasting models, while they focused on zero-shot using pre-trained language models. However, none of them have investigated the importance of cross-attention for time series forecasting.
@@
Fixed temporal order in time series is the distinct property of time series, in contrast to the language domain where semantic information does not heavily depend on the word ordering [4]. Thus, some researchers have used learnable positional encoding in Transformers to embed time-dependent properties [9, 23]. However, Zeng et al. [26] first argued that self-attention is not suitable for time series due to its permutation invariant and anti-order properties. While they focus on building complex representations, they are inefficient in maintaining the original context of historical and future values. They rather proposed linear models without any embedding layer and demonstrated that it can achieve better performance than Transformer models, particularly showing robust performance to long input sequences. Recent linear time series models outperformed previous Transformer models with simple architectures by focusing on pre-processing and frequency-based properties [10, 2, 21]. On the other hand, Woo et al. [22] investigated the new line of works of time-index models, which try to model the underlying dynamics with given time stamps. These related works imply that preserving the order of time series sequences plays a crucial role in time series forecasting.
3 Revisiting Self-Attention in Time Series Forecasting
Motivation of Self-Attention Removal
Following the concerns about the effectiveness of self-attention on temporal information preservation [26], we conduct an experiment using PatchTST [14]. We consider three variations of the PatchTST model: the original PatchTST with overlapping patches with length 16 and stride 8 (Fig. 2a); a modified PatchTST with non-overlapping patches with length 24 (Fig. 2b); and a version where self-attention is replaced by a linear embedding layer, using non-overlapping patches with length 24 (Fig. 2c). This setup allows us to isolate the effects of self-attention on temporal information preservation, while controlling for the impact of patch overlap.
(a) Original PatchTST
(b) PatchTST w/ non-overlapping
(c) PatchTST w/o self-attn
Figure 2: Absolute values of weights in the final linear layer for different PatchTST variations. The distinct patterns reveal how each model captures temporal information.
Fig. 2 illustrates the absolute values of the weights in the final linear layer for these model variations. Compared to the original PatchTST (Fig. 2a), both non-overlapping versions (Fig. 2b and Fig. 2c) show more vivid patterns. The version with linear embedding (Fig. 2c) demonstrates the clearest capture of temporal information, suggesting that the self-attention mechanism itself may not be necessary for capturing temporal information.
In Table 3, we summarize the forecasting performance of the original PatchTST (Fig. 2a) and PatchTST without self-attention (Fig. 2c). PatchTST without self-attention consistently improves or maintains performance across all forecasting horizons. Specifically, the original version with self-attention shows lower performance for longer forecast horizons. This result suggests that self-attention may not only be unnecessary for effective time series forecasting but could even hinder
[…]
tableEffect of self-attention in PatchTST on forecasting performance (MSE) on ETTm1. Horizon original w/o self-attn 96 0.290 0.290 192 0.332 0.328 336 0.366 0.359 720 0.416 0.414
performance in certain cases. Therefore, better performance of w/o self-attention challenges the conventional belief regarding the importance of self-attention mechanisms in Transformer-based models for time series forecasting tasks.
Our findings offer new insights into the role of self-attention in time series forecasting. As shown in Fig. 2 and Table 3, replacing self-attention with a linear layer not only captures clear temporal patterns but also results in significant performance improvements, particularly for longer forecast horizons. These results highlight potential areas for enhancing the handling of temporal information, beyond addressing the well-known concerns regarding computational complexity.
Rethinking Transformer Design
@@ 3 Revisiting Self-Attention in Time Series Forecasting
Given the challenges associated with self-attention in time series forecasting, we propose a fundamental rethinking of the Transformer architecture for this task. Fig. 3 illustrates the differences between existing architectures and our proposed approach. Traditional Transformer architectures (Fig. 3a) and encoder-only models (Fig. 3b) rely heavily on self-attention mechanisms, which may lead to temporal information loss. In contrast, Zeng et al. [26] proposed a simplified linear model, DLinear (Fig. 3c), which removes all Transformer-based components. While this approach reduces computational load and potentially avoids some temporal information loss, it may struggle to capture complex temporal dependencies.
To address these challenges while preserving the advantages of Transformer architectures, we propose the Cross-Attention-only Time Series transformer (CATS), depicted in Fig. 3d. Our approach removes all self-attention layers and focuses solely on cross-attention, aiming to better capture temporal dependencies while maintaining the structural advantages of the transformer architecture. In the following section, we will introduce our CATS model in detail, explaining our key innovations including a novel use of cross-attention, efficient parameter sharing, and adaptive masking techniques.
4 Proposed Methodology
@@ 4.1 Problem Definition and Notations
In traditional time series transformers, we feed the historical multivariate time series $\bm{X}$ to embedding layers, resulting in the historical embedding $\bm{H}\in\mathbb{R}^{D\times L}$. Here, $D$ is the embedding size. Note that, in channel-independence manners, the multivariate input is considered to separate univariate time series ${\mathbf{x}}\in\mathbb{R}^{1\times L}$. With patching [14], univariate time series ${\mathbf{x}}$ transforms into patches ${\mathbf{p}}=\texttt{Patch}({\mathbf{x}})\in\mathbb{R}^{P\times N_{L}}$ where $P$ is the size of each patch and $N_{L}$ is the number of input patches. Similar to non-patching situations, patches are fed to embedding layers $\bm{P}=\texttt{Embedding}({\mathbf{p}})\in\mathbb{R}^{D\times N_{L}}$.
@@ 4.2 Model Structure
Our proposed architecture consists of three key components: (A) Cross-Attention with Future as Query, (B) Parameter Sharing across Horizons, and (C) Query-Adaptive Masking. Fig. 4 illustrates how our model modifies the traditional transformer structure.
By focusing solely on cross-attention, our approach allows us to maintain the periodic properties of time series, which self-attention, with its permutation-invariant and anti-order characteristics, struggles to capture. Furthermore, we leverage advanced architectural designs of time series transformers, such as patching [14], which linear models cannot utilize. The following subsections provide detailed descriptions of each component of our CATS model, explaining how these elements work together to address the challenges of time series forecasting.
@@
Similar to self-attention, the cross-attention mechanism employs three elements: key, query, and value. The distinctive feature of cross-attention is that the query originates from a different source than the key or value. Generally, the query component aims to identify the most relevant information among the keys and uses it to extract crucial data from the values [1, 27]. In the realm of time series forecasting, where predictions are often made for a specific target horizon—such as forecasting 10 steps ahead. Therefore, within this concept of forecasting, we argue that each future horizon should be regarded as a question, i.e., an independent query.
To implement this, we establish horizon-dependent parameters as learnable queries. As shown in Fig. 4, we begin by creating parameters for the specified forecasting horizon. For each of these virtualized parameters, we assign a fixed number of parameters to represent the corresponding horizon as learnable queries ${\mathbf{q}}$. For example, ${\mathbf{q}}_{i}$ is a horizon-dependent query at $L+i$. When patching is applied, these queries are then processed independently; each learnable query ${\mathbf{q}}\in\mathbb{R}^{P}$ is first fed into the embedding layer, and then fed into the multi-head attention with the embedded input time series patches as the key and value.
@@
Transformer [13] $\mathcal{O}(L^{2})$ $\mathcal{O}(T(T+L))$ 10.4ms Informer [29] $\mathcal{O}(L\log L)$ $\mathcal{O}(T(T+\log L))$ 13.5ms
Autoformer [23] $\mathcal{O}(L\log L)$ $\mathcal{O}\left(\left(L/2+H\right)\log\left(L/2+T\right)\right)$ 24.1ms Pyraformer [11] $\mathcal{O}(L)$ $\mathcal{O}(T(T+L))$ 11.2ms
FEDformer [31] $\mathcal{O}(L)$ $\mathcal{O}\left(L/2+H\right)$ 69.3ms Crossformer [28] $\mathcal{O}\left({M}L^{2}/{P^{2}}\right)$ $\mathcal{O}\left({M}T(T+L)/{P^{2}}\right)$ 30.6ms
PatchTST [14] $\mathcal{O}\left({L^{2}}/{P^{2}}\right)$ 7.6ms CATS (Ours) - $\mathcal{O}\left({LT}/{P^{2}}\right)$ 7.0ms
@@
This independent forecasting mechanism offers a notable advantage; a higher level of parameter sharing. As demonstrated in [14], significant reductions in the required number of parameters can be achieved in time series forecasting through parameter sharing between inputs or patches, enhancing computational efficiency. Regarding this, we propose parameter sharing across all possible layers — the embedding layer, multi-head attention, and projection layer — for every horizon-dependent query ${\mathbf{q}}$. In other words, all horizon queries ${\mathbf{q}}_{1},\dots,{\mathbf{q}}_{T}$ or ${\mathbf{q}}_{1},\dots,{\mathbf{q}}_{{N_{T}}}$ share the same embedding layer used for the input time series ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{L}$ or patches ${\mathbf{p}}_{1},\dots,{\mathbf{p}}_{N_{L}}$ before proceeding to the cross-attention layer, respectively. Furthermore, to maximize the parameter sharing, we also propose cross-dimension sharing that uses the same query parameters for all dimensions.
For the multi-head attention and projection layers, we apply the same algorithm across horizons. Notably, unlike the approach in PatchTST [14], we also share the projection layer for each prediction. Specifically, PatchTST, being an encoder-only model, employs a fully connected layer as the projection layer for the encoder outputs $\bm{P}\in\mathbb{R}^{D\times N_{L}}$, resulting in $(D\times N_{L})\times T$ parameters. In contrast, our model first processes raw queries ${\mathbf{q}}=[{\mathbf{q}}_{1},\ldots,{\mathbf{q}}_{N_{T}}]\in\mathbb{R}^{P\times N_{T}}$. These queries are then embedded through the cross-attention mechanism, resulting in $\bm{Q}=[\bm{q}_{1},\ldots,\bm{q}_{N_{T}}]\in\mathbb{R}^{D\times N_{T}}$. The final projection uses shared parameters $W\in\mathbb{R}^{P\times D}$, producing an output $W\bm{Q}\in\mathbb{R}^{P\times N_{T}}$. Thus, our number of parameters for this projection becomes $P\times D$, which is not proportionally increasing to $T$. This approach significantly reduces time complexity during both the training and inference phases.
@@
Therefore, to ensure the model focuses on each horizon-dependent query ${\mathbf{q}}$, we introduce a new technique that masks the attention outputs. As illustrated in the right-bottom figure of Fig. 4, for each horizon, we apply a mask to the direct connection from Multi-Head Attention to LayerNorm with a probability $p$. This mask prevents access to the input time series information, resulting in only the query to influence future value predictions. This selective disconnection, rather than the application of dropout in the residual connections, helps the layers to concentrate more effectively on the forecasting queries. We note that this approach can be related to stochastic depth in residual networks [8]. The stochastic depth technique has proven effective across various tasks, such as vision tasks [17, 25]. To the best of our knowledge, this is the first application of stochastic depth in Transformers for time series forecasting. A detailed analysis of query-adaptive masking can be found in Appendix.